OpenStudy (anonymous):

find the vertex,line of symetry, and minimum maximum value of the equation f(x)=2x^2+2x+3

OpenStudy (anonymous):

get first derivative

OpenStudy (anonymous):

i have x = -1/2 y=5/2 line of symetry =5/2 and minimum =-1/2

OpenStudy (anonymous):

get first derivative

OpenStudy (anonymous):

and how did you go about findingthat

OpenStudy (anonymous):

-b/2a=-2/4=-1/2 this is the vertex f(-1/2)=2(-1/2)^2+3=5/2 line of symetry =5/2 and minimum =-1/2

OpenStudy (dumbcow):

hey awesome, good job however, you need to switch your line of symmetry and minimum max/min refer to the y-value remember line of symmetry is that vertical line that cuts through the vertex

OpenStudy (dumbcow):

@ mathcruncher, this is not for a calculus class

OpenStudy (anonymous):

you can also do it by completing the square =2(x^2+x+3/2) =2((x+0.5)^2-1/4+3/2) =2((x+0.5)^2+5/4) =2((x+0.5)^2)+5/2 line of symmetry will be equating (x+0.5)=0 therefore x=-0.5 x coor of vertex will be -0.5and y-coor will be subs x=-0.5 in the equation which is 5/2 it is a minimum bcoz the coeff of x^2 is positive, min at x=-0.5 adn min value will be -5/2